LjTheBaller16:

A car seller buys a car from a manufacturer for $8,000. He increases the cost by 6 percent. What is the markup value?

darkknight:

find 6% of 8000$ then add that to $8000

XxAngelGirlxX:

Mark up value is either a fixed amount or a percentage of the total cost or selling price. In this problem, mark up value is the percentage of the total cost. To determine the retail price, total cost and mark up must be added. Selling Price = Total Cost + Mark up value based on Total Cost However, we are looking for the Total cost. Thus, our formula should be Total Cost = Selling Price - Mark up value based on Total Cost. Let X = total cost. Selling Price = $8,000 Mark up vale = 6% of total cost. X = $8,000 - 6%X X = $8,000 - 0.06X To get X, transfer -0.06X to the other side and change its sign from negative to positive. X + 0.06X = $8000 1.06X = $8000 To get X, divide both sides by 1.06 1.06X / 1.06 = $8000 / 1.06 X = $7,547.17 total cost. The problem is looking for the mark up value and since it states that the mark up value is 6% of the total cost, then: Total Cost x 6% = Mark up value $7,547.17 x 0.06 = $452.83 mark up value To check: X + 0.06X = $8000 $7547.17 + $452.83 = $8000 $8000 = $8000 equal. The answer is $480 8,000 x 6% = 480

darkknight:

how did u type all that up so fast, copy paste?

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight how did u type all that up so fast, copy paste? \(\color{#0cbb34}{\text{End of Quote}}\) Shh, she's a super typer

LjTheBaller16:

lmao

LjTheBaller16:

thx i got a 100 on my quiz

darkknight:

please don't post quiz questions next time. Against our policy

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight please don't post quiz questions next time. Against our policy \(\color{#0cbb34}{\text{End of Quote}}\) wym our policy are u the owner of the site

darkknight:

just a member of the community : ) but we can't help with test or quiz questions

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight just a member of the community : ) but we can't help with test or quiz questions \(\color{#0cbb34}{\text{End of Quote}}\) yea okay sir or ma'am idc